OpenStudy (anonymous):

a random variable R is normally distributed with a mean of 7 and a standard deviation of 1.5 . Find the value of W so that P (8.8

OpenStudy (mathmate):

Z-value for 8.8=(8.8-7)/1.5=1.2 Look up for what value of Z=w such that P(R<w)= P(R<1.2) +0.08.